OpenStudy (anonymous):

Which of the following expressions is not equivalent to 1?

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Where are the choices?

OpenStudy (anonymous):

x^0, (ab)^0, 2p^0, 1/w^0

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Anything to the power of 0 is 1..

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so it is D?

OpenStudy (igreen):

No..

OpenStudy (igreen):

1 / w^0 let's say w = 2 1 / 2^0 1 / 1 1

OpenStudy (anonymous):

uhh

OpenStudy (anonymous):

oh okay

OpenStudy (igreen):

C should be: \(2 \times p^0\)

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so it is A

OpenStudy (igreen):

No.. x^0 2^0 1

OpenStudy (igreen):

(ab)^0 (2)(3)^0 What does this equal?

OpenStudy (anonymous):

0

OpenStudy (igreen):

No..anything to the power of 0 is 1.

OpenStudy (igreen):

2 * 3 is 6 so we have 6^0 And what is that?

OpenStudy (anonymous):

6

OpenStudy (igreen):

No..I just said anything to the power of 0 is 1.

OpenStudy (anonymous):

o 1

OpenStudy (anonymous):

it is 1

OpenStudy (igreen):

Yes.. So that's incorrect.

OpenStudy (anonymous):

so it is A right?

OpenStudy (igreen):

Now let's look at option C: \(2 \times p^0\) Let's say 'p' is 2: \(2 \times 2^0\) What does 2^0 equal?

OpenStudy (anonymous):

1

OpenStudy (igreen):

Yes.. that gives us: \(2 \times 1\) And what does that equal?

OpenStudy (igreen):

@shifuyanli

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@iGreen?

OpenStudy (igreen):

Yes..so what will your answer be since C does NOT equal 1?

OpenStudy (anonymous):

um C

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thanks